OpenStudy (anonymous):

What about this inequality how would it be solved? sqrt[3]{4m+1}-2=2

OpenStudy (anonymous):

That's an equality, not an inequality. \[ \sqrt[3]{4m+1} - 2 = 2\] \[ \sqrt[3]{4m+1} = 4\] \[4m+1 = 4^3 = 64\] \[4m = 63\] \[m = \frac{63}{4} \]

OpenStudy (anonymous):

ohh okay i get it thanks !